# An Industry Interview Study of Software Signing for Supply Chain Security

### v1

@@ 5.3.2 As a Secondary Security technique
Some subjects believe that software signing is a supplementary method used to ensure the integrity of primary security techniques. They often deemed it important only in conjunction with methods like Attestations and other Manifest forms such as the SBOM. As S2 said, “I think the most important thing is accurate metadata collection and centralization. And even if that data’s not signed, we can still get a lot of usefulness out of it. Signatures, in my opinion, they offer an additional layer of security on top of what we should be doing, which is metadata collection”. S17 similarly views signing as a final check on top of other security practices: “Signing was essentially the way of saying, ‘I have checked, we did all the [security practices]’.”.
5.4 RQ4: Impact of Supply Chain Security Awareness on Software Signing Practices
Finding 4. 1. Software supply chain failures usually require immediate fixes to the malfunctioning or compromised components. Large-scale failures typically trigger a comprehensive reassessment, leading to changes in the security strategy of a team or organization. 2. Security regulations, frameworks, and standards drive the adoption of signing and other security techniques, but some subjects criticized these obligations.
@@ 5.4 RQ4: Impact of Supply Chain Security Awareness on Software Signing Practices
In this section, we explore how experience with software supply chain security influences a company’s decision to adopt and implement software signing. We asked subjects about their personal and organizational experiences with software supply chain failures (attacks, vulnerabilities, and other incidents) and how these experiences affected their team’s or organization’s software supply chain security decisions, particularly regarding software signing. Additionally, we queried subjects on the impact of software supply chain security frameworks, standards, and regulations on their use of software signing and other security techniques.
5.4.1 How experiences with SSC attacks influence the use of software signing
@@
TABLE V: Experienced software supply chain failures and type of Security change triggered.
@@
Type # Subjects (# Orgs) # Direct fix # $\Delta$Signing # $\Delta$General Security process
Attack 4 (4) 4 2
Vulnerability 9 (6) 9 1 2
Other Incidents 5 (5) 5 2
Software supply chain security failures typically have direct fixes, except in cases of severe attacks. While only 4 of our subjects indicated experiencing software supply chain attacks, only one of these attacks (where the subject was hired to help fix the aftermath) was highlighted as major, resulting in collateral damage and leading to a complete overhaul of the software and security processes. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S4: “I would say that at this point it’s probably difficult to find any process at that organization that wasn’t inspected and potentially changed as a result of that. Every aspect of how everything was built there was changed … There was no choice in that situation.”
Some subjects report that software supply chain security failures led to a broader change in their team’s or organization’s security posture. For example, S18 describes their team’s response to the Log4j vulnerability, which was present in their system. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S18: “… it was a reality check to know that even companies like ours … can be attacked through open source libraries, … Asset inventory is the biggest thing that is lacking where you don’t know what is currently running in your environment. And if it is running, you don’t know how old that is, or what are the dependencies which is currently running in production. So that pushed us towards making sure that we have a continuous inventory of container images, which are currently running, for example, for Kubernetes, and then there are initiatives like GUAC.”
Generally, most software supply chain reported by our subjects were not severe, and as such most of them received direct fixes. This could be swapping out a defaulting tool or dependency, re-releasing an affected release, publishing security advisories, etc. We highlight some of those responses. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S5: “So what happens is that, if the current version where that security risk is present, if that is also publicly available, we have to create a security advisory to let everyone know that if you are working with this version of our software, you should upgrade to the next version which contains the fix. And we let them know what the fix is and if they are affected by the fix.” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S1: “to my knowledge it did not cause a major restructuring or change that is … Yeah, I would say that they probably patched any of the security issues that they had.”
The only instance where a signing-related fix directly resulted from a software supply chain security failure was when the subject team’s product was a safety-critical application. This team added intoto’s signed attestation capability. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S2: “we started reaching for the In-toto project, and we saw that it allowed you to decouple the security metadata collection and validation. You can abstract that away from the actual build process itself. So that decoupling allowed us to solve this problem of how do we know that this opaque process for the software, that opaque certification, how do we know that was followed without actually knowing the details of what that process is?”
5.4.2 Influence of security regulations, frameworks, and standards
TABLE VI: Regulatory and Standard Influences on Security Technique Adoption: Analysis of Subjects’ Responses. 4 Subjects say they are not influenced by either.
ANY INFLUENCE?
#Subjects (Yes) #Subjects (None)
Regulation(EO-14028) 11 7
Other Standards 8 10
The adoption of software signing and other security techniques is heavily influenced by regulations and other standards. Specifically, we asked our subjects, “Are these strategies (software signing implementation & other complementary security methods(earlier listed by subjects)) influenced by regulations and standards?”. While answering generally, most of our subjects agreed that regulations, particularly the E0-14028, played a key role in the way their team and organization have implemented software supply chain security. Although software signing is not explicitly mentioned, Some subjects did note that their use of SBOMs was directly a result of regulations or standards. We only had 1 subject who noted that the standards influenced their use of signing in certifying attestations(SBOMs).
@@ 5.4.2 Influence of security regulations, frameworks, and standards
S1 highlights this influence of the executive order: “Yeah, I would say certainly the collection of SBOMs and doing that more broadly has been a response to things like the executive order, at least for our team specifically. I would say, yeah, the executive order played just a strong role in what we are starting to do now. Whether or not some teams are now trying to be preventative, that’s a good question.” S7 also adds to S1’s point: “we keep the executive orders in mind when we build our product…”
Some of our subjects (3) only relied on standards (not on regulations) for security recommendations. For example: {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S18: “… as an organization, currently we don’t have any government contracts. And I think the executive order is mostly targeted towards companies, organizations who are having contracts with the government and they need to produce certain SBOMs and things like those. So our push towards supply chain or better security was not influenced by the executive order.”
S18 also adds: “But I think in terms of where we are at, I think in terms of frameworks, as you mentioned, we do look at some of these frameworks like SLSA,… There are quite a few …, and each of them target different parts of the supply chain. But as an organization, it’s really difficult to keep track and follow each of these standards of frameworks, to be honest, because we have limited resources, we can’t just keep on handling those. So we are looking at SLSA, or at least trying to get closer to where we should be for the SLSA level. Again, it’s not written in stone that we need to have level one, level two, level three, things like that, but that’s something which we strive to achieve sooner.”
S14 adds to this line of thought: “The recent interests from the US federal government and software supply chain security, SBOMs, and the executive order from 2021, it’s hard to detect very much influence on our security because of those … I think when observers of software security read that executive order and then commentary in the aftermath, they often think that SBOMs have become mandatory…”
5.4.3 Criticism of Regulations and Standards
Also, our subjects identified several issues with the current set of software supply chain security standards and regulations that hinder their influence on the adoption of software signing and other security tools. We highlight four kinds of issues.
(a) Regulations don’t specify standard data exchange format for security tools: Criticism of standards and regulations here lies in their inadequacies to specify or promote standard data interchange formats. This is highlighted by the lack of data formats for certain current security techniques such as Software Bill of Materials (SBOMs), other manifest-type techniques, signing tools, etc. The absence of standards for these techniques leads to numerous undesired practices, such as the use of vulnerability scan reports as an input for SBOM data. This ultimately leads to compatibility problems among different security techniques.
{quoting}
[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S16: “…one of the gaps with the executive order is, yeah, you give an SBOM now. But what are they going to do with it? Nobody knows yet. And how do you compare it? Because even right now, going back to the tools, there are different versions of SPDX. There are different versions of Cyclone DX. They’re the two big ones. This goes back to tools. There are a bunch of tools out there that will convert say 2.0 back to three and you come up with different answers using different tools.” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S3: “SBOMs are a good example of something that has a standard, but all of the implementations end up being different enough that they’re very difficult to translate from one format to the other or even get consistency within the same format type.”
(b) Regulation and standards are just checkbox rules:
Here, subjects argue that standards and regulations function more as compliance checks rather than as security enhancers. They point out that these standards and regulations lack security enhancement features, such as implementation models. Instead, they primarily promote compliance-type features, such as guidelines and best practices lists. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S17: “So there is currently a disconnect between regulations and security. So compliance and regulations keep people out of jail, keep people from getting fined, [while] security stops you from getting hacked. And the problem there is sometimes those two things can conflict. And compliance is there to make sure that everybody is following reasonable sets of rules and security is there to actually implement that and create the software and the processes that hit that … Regulations help provide the requirement. They don’t necessarily make security better on their own. And there’s always going to be a balance because what compliance tells you is compliance helps you with making sure that everybody’s following the standard … the only way for me to be compliant to this is to ignore the actual problem and just check the box.”
(c) Difficult to implement in the SDLC: A group of participants highlighted the challenges encountered when attempting to integrate the recommendations of security standards into their software engineering processes. These challenges primarily arise from the difficulties in training engineers on the application of these standards, and the deficiencies in the technical implementation of the standards themselves. S12 makes this point clear in their response; {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S12: “… just having the executive order was just another way to explain it to the higher-ups… I don’t think that the executive order in particular affected the business … There’s kind of milestones. You don’t need to know exactly what the ins and outs are technically of achieving them … But it’s a lot of work getting that in place and upskilling all of your existing developers, et cetera. So yeah, anything that requires going to all 4,000, 10,000 of your developers in some of these massive organizations and making sure every single one of them does something different, it is a lot of work.”
(d) Standards are not formalized: Criticism regarding the formalization of standards, as expressed by some participants, reflects the lack of consensus surrounding these standards. Participants argue that since these standards are not necessarily created by a majority of industry consortia, they should not be referred to as ‘standards’. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S13: “the term standards is imprecise. So some people think of standards as being very formalized but some people would say any documented practice is a standard. So there’s the SLSA project in it, and one might argue that that is a standard or one might argue that it’s not a standard, but it is a documented set of recommendations. And if people are implementing those, you could argue that they’re implementing a standard.”
@@ 6.1 Recommendations from this study
Improving Signature & Artifacts Verification in Signing workflows.   Our refined software supply chain factory model analysis has highlighted a significant gap in the current practice of signing - signature verification. This issue was underscored by several subjects in our study.
@@
Further, while artifact verification, in this context, refers to conducting a thorough check of the integrity of the software artifact by means of identifying the signer. However, the use of signed attestations (i.e., are statements about a particular action or property in the supply chain) can further verify the authenticity and integrity of the process and will become crucial to further improve the security posture of organizations.
Future works in this area should focus on proposing improved tooling around existing or new software signing tools to include these capabilities in their design. This enhancement would not only improve transparency but also reduce the costs and expertise requirements for setting up such infrastructure.
By addressing these issues, we can ensure that signing is not just a procedural element, and instead plays a crucial role in maintaining the integrity and authenticity of the software development process.
Addressing Incentive Challenges in Software Signing.  One of the most frequently reported challenges in the realm of software signing is the lack of proper incentives for its adoption. These incentives can range from support from management, the ease of using signing tools, as well as comprehensive education on the importance and application of signing.
While some of these issues, such as the need for educating practitioners on the significance and application of signing, can be addressed with relative ease, others necessitate more elaborate planning. For instance, fostering management support might require demonstrating the tangible benefits of software signing in terms of enhanced security and reduced risk of software tampering.
Moreover, though the user-friendlness of signing software has been evidenced by various “why Johnny can’t” papers, more research is needed in lowering the adoption bar for software signing systems.
Regardless of the complexity of these challenges, investing in these areas is paramount for enhancing an organization’s security posture. By fostering a culture that values and understands the importance of software signing, organizations can ensure the integrity of their software and protect against potential security threats.
@@
Criticism of Current Standards and Regulations.  A significant finding of our study is the criticism directed towards current standards and regulations, particularly concerning their creation process and the difficulty of their implementation. These issues could be mitigated by ensuring that these standards and regulations are informed by a broader section of the industry consortium. This would involve gathering input from a diverse range of practitioners, including those from different sectors, roles, and levels of expertise.
While frameworks like SLSA already involve collaboration among multiple organizations, we advocate for a more inclusive effort. This would target a larger group of practitioners, ensuring that the standards and regulations are not only comprehensive but also practical and implementable across various contexts.
In the future, research could focus on exploring effective strategies for fostering such inclusive collaborations and assessing their impact on the quality and applicability of the resulting standards and regulations. By addressing these issues, we can work towards standards and regulations that are not only robust and comprehensive but also widely accepted and easily implementable. This would significantly enhance the security and integrity of software signing practices across the industry.
@@ 6.2 Contrast to prior work
Distinctions of Refined Model Compared to Existing Frameworks.  While all software supply chain security frameworks agree on the defined threat model, our refinement of this model differs from existing frameworks in several ways.
@@ 7 Conclusion
Signature-based supply chain provenance is a topic of increasing importance, driven by cyberattacks, executive orders, and new industry standards. Although the literature contains many studies of signing as a technology, there have been few studies of the use of signing in engineering contexts. Those that exist have focused on open-source, or on the mechanics of signing. In this paper, we present the first qualitative examination of signing in engineering organizations’ security strategies. We interviewed 18 high-ranking practitioners across 13 organizations, including C-suite staff, principal and senior engineers, managers, and consultants. Our four main findings are that (1) Software products are being signed, but the signatures are not yet being validated; (2) The factors that hamper signing include infrastructure and compatibility issues, lack of expertise, developer and organizational attitudes, and lack of resources; (3) Signing is considered an important guarantee of provenance, but it is not the main technique these organizations consider in their cybersecurity postures; and (4) Organizational behaviors are being influenced by major failures (e.g., SolarWinds) and by regulations and industry standards, but some of our subjects expressed concern about the development process and quality of those recommendations. Our results indicate many opportunities for technical and operational advances in the use of software signing to improve software supply chain provenance.
8 Replicability
For anonymity reasons, we do not plan to share the raw transcripts of our interviews. However, our interview protocol, codebook, etc. are available in the Appendices.

### v2

@@ 5.3.2 Signing as a Secondary Security Technique
Some subjects believe that software signing is a supplementary method used to ensure the integrity of primary security techniques. 7 subjects (6 organizations) had the opinion that Software signing does not provide enough guarantee for provenance. They often deemed it important only in conjunction with methods like Attestations and other Manifest forms such as the SBOM. For these organizations, signing is a valuable part of a broader, multi-layered security strategy. Some subjects who view signing as a strong guarantee of provenance also emphasized that adding a manifest or attestation further strengthens this guarantee. As S2 (SAAS, SSC Security tool) said, “I think the most important thing is accurate metadata collection and centralization. And even if that data’s not signed, we can still get a lot of usefulness out of it. Signatures, in my opinion, they offer an additional layer of security on top of what we should be doing, which is metadata collection.” S17 (SAAS, SSC Security tool) similarly views signing as a final check on top of other security practices: “Signing was essentially the way of saying, ‘I have checked, we did all the [security practices]’.”
5.3.3 As a Requirement-Driven Practice
Subjects in this category indicated that software signing was primarily used as a "check the box" technique to satisfy regulatory, framework, customer or organizational requirements. They did not view signing as a primary security measure. Instead, it was employed mainly to meet regulatory or contractual obligations, rather than being treated as an essential part of their security strategy. For these subjects, signing was often seen as a procedural necessity—for instance, to submit a pull request, initiate a build, or make a commit—rather than as a measure to enhance security. 4 subjects representing 3 organizations fall into this category. {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S17 (SAAS, SSC Security tool): “… at certain organizations it felt like it (software signing) was more of a check-the-box thing of like, "Hey, we should sign it because that’s what people tell us to do” {quoting}[leftmargin=0.15cm,rightmargin=0.15cm,font=italic] S8 (Internal security tool & cloud APIs, Internet services): “…for signed binaries and code signing …a lot of those [requirements] are driven…by platform requirements…to prevent us from being blocked from…these platforms, code signing…was largely driven by [these platforms].”
5.4 RQ4: Impact of Internal & External Events on Software Signing Practices
In this section, we explore the impact of two kinds of security events: security failures such as Log4j, and the advent of security standards and regulations:
  1. 1.
    Experienced Software Supply Chain Failures: We consider these internal if a subject’s organization is a direct victim of a Software Supply Chain attack, and external if a subject’s organization is not directly impacted in a Software Supply Chain event, e.g., a general Software Supply Chain incident like the Solarwinds attack.
  2. 2.
    Software Supply Chain Standards: These are industry wide external events e.g., the publishing of the EO-14028 executive order on Software Supply Chain security.
@@ 5.4 RQ4: Impact of Internal & External Events on Software Signing Practices
We asked subjects about their organizational experiences with software supply chain failures (attacks, vulnerabilities, and other incidents) and how these experiences affected their team’s or organization’s software supply chain security decisions, particularly regarding software signing. Additionally, we queried subjects on the impact of software supply chain security frameworks, standards, and regulations on their use of software signing and other security techniques.
In their responses, subjects described a security failure they had experienced. We categorized each failure as a Vulnerability (unexploited security flaws), an Incident (SSC failures with no known malicious intent), or an Attack (deliberate and malicious failures). Then, subjects described changes these failures prompted in the organization. We categorized these as Direct Fixes (patched but no change to organizational security approach), General Security Process (large-scale changes involving multiple security techniques), and Signing (adoption or change in the implementation of software signing).
5.4.1 Impact of Experiences with SSC Failures
@@
Table 5: Kinds of software supply chain failures experienced, and associated security changes.
@@
Type # Subjects (# Orgs) # Direct fix # Change in General Security process # Change in Signing process
Vulnerability 9 (6) 9 2 1
Incident 5 (5) 5 2 0
Attack 4 (4) 4 2 0
Table 5 summarizes these events and the resulting changes. Most cases were non-malicious, either Vulnerabilities or Incidents. However, four subjects had experienced attacks, with one major attack leading to extensive collateral damage and a complete overhaul of software and security processes.
Across these events, the typical resolution was via Direct Fix, without an infrastructural change in the general security process or the signing process specifically. When they occurred, changes to the general security process typically involved adoption of a group of tooling such as for SBOMs. As a result of one event, a vulnerability led to changes in the signing process. In this case, the team (which had a safety-critical application) integrated in-toto’s signed attestation capability. The subject S2 (SAAS, SSC Security tool) stated: “We saw that [in-toto] allowed you to decouple the security metadata collection and validation…that allowed us to solve this problem of how do we know that this opaque process for the software, that opaque certification…was followed without…knowing the details of…that process.”
5.4.2 Impact of Security Regulations, Frameworks, and Standards
Our subjects reported that security frameworks, standards, and regulations had no direct impact on their organizations’ adoption of software signing as a security method. However, these standards and regulations did influence the general adoption of other security techniques, such as SBOM. 9 participants from 5 organizations shared this view, while only 2 (from 2 organizations) from this group explicitly linked this to signing. 5 participants (from 4 organizations) provided unclear statements (personal opinions or unfamiliarity with how decisions are made).
@@ 5.4.2 Impact of Security Regulations, Frameworks, and Standards
We asked our subjects, “Are [your SSC security] strategies influenced by regulations and standards?”. S1 (POC software, Digital technology)’s response was typical: “Yeah, I would say certainly the collection of SBOMs and doing that more broadly has been a response to things like the executive order, at least for our team specifically…The executive order played just a strong role in what we are starting to do now. Whether some teams are now trying to be preventative, that’s a good question.”
@@ 6 Discussion
6.1 Impact of Organizational Size and Type of Produced Software on Signing Practices
We did not observe any significant effect of organizational size (and nature of software product) on where signing was implemented (RQ1,  Figure 3). However, we did note patterns in the challenges reported (RQ2) and the influence of standards and regulations on the adoption of software signing (RQ4). It is important to note that these observations represent correlations rather than causations. While our findings highlight potential relationships, further study is necessary to establish causative factors and better understand the mechanisms driving these patterns.
Effect on Challenges.  We observed a noticeable difference in the organization size of participants who reported certain type of challenges. Only subjects from large organizations (S6, S8, S13, S15) talked about difficulties in operationalizing the signing process. We conjecture this is because of the greater coordination effort required to harmonize the engineering process of several teams.
The nature of the software product also seem to exert some influence on the challenge experienced. Unsurprisingly, only subjects from organizations with non-security-focused product areas (S11 (Firmware & testing software, Social technology) and S13 (Internal security tool, Telecommunication)) reported a lack of management incentive as a barrier to signing. While Only participants from security-focused product teams (S3 (SAAS, SSC Security tool), S10 (SAAS, Dev tools)) reported challenges in creating effective signing policies.
Effect on Influence of Standards and Regulations.  As we reported, nine participants from five organizations noted that standards and regulations influenced the adoption of tools like SBOM. Within this group, only two participants (from two of the five organizations) explicitly linked this to signing. Notably, all five organizations were either large or focused on security-related products.
6.2 Relation between Participants’ perceived importance of Signing and Signing Implementation
We remark on a recurring dissonance between the perceived importance of Software signing by the participants and their implementation practices. Although all organizations signed their final products, signatures were rarely used to verify integrity or provenance Table 3. Ten subjects (from nine organizations) regarded Software signing as a very strong guarantee of provenance, but only three participants (from two organizations) reported verifying signatures for internal contributors, and none did so for external contributors. This suggests a unidirectional use of Software signing, focusing on establishing trust for their outputs while placing lower expectations on the provenance of their inputs. As a result, these organizations are effectively publishing software with a higher degree of provenance than they demand from their dependencies or external contributions.
@@ 6.3 Recommendations from this study
Improving Signature & Artifacts Verification in Signing Workflows.   Our refined software supply chain factory model has highlighted a significant gap in the current practice of signature verification. This issue was underscored by several subjects in our study.
@@
Effective artifact verification is essential to ensure the integrity of the software. In this context, it involves a thorough check of the integrity of the artifact by identifying the signer. Moreover, the use of signed attestations—statements verifying specific actions or properties within the supply chain—can further authenticate and safeguard the process. Adopting these practices will be crucial for enhancing the security posture of organizations.
Future works in this area should focus on proposing improved tooling around existing or new software signing tools to include these capabilities in their design. This enhancement would not only improve transparency but also reduce the costs and expertise requirements for setting up such infrastructure. Also, while much of our recommendations focus on “improving tooling”, further research is needed to define what kinds of tooling would qualify. This research might involve experimental and empirical analyses of the current generation of software signing tools, particularly in terms of functionality, integration [65], and usability.
As additional takeaways, we emphasize the need for further research to address critical trust gaps within the open-source ecosystem. Our findings ( section 5.1.4) highlight an underlying lack of trust in open-source packages, suggesting that cross-ecosystem provenance maintenance could serve as a potential solution.
By addressing these issues, we can ensure that signing is not just a procedural element and instead plays a crucial role in maintaining the integrity and authenticity of the software development process.
Addressing Incentive Challenges in Software Signing.  One of the most frequently reported challenges in the realm of software signing is the lack of proper incentives for its adoption. These incentives can range from support from management, the ease of using signing tools, as well as comprehensive education on the importance and application of signing. Although some of these issues, such as the need for educating practitioners on the significance and application of signing, can be addressed with relative ease, others necessitate more elaborate planning. For instance, fostering management support might require demonstrating the tangible benefits of software signing in terms of enhanced security and reduced risk of software tampering. Moreover, though the user-friendliness of signing software has been evidenced by various “Why Johnny Can’t” papers, more research is needed in lowering the adoption bar for software signing systems. Regardless of the complexity of these challenges, investing in these areas is paramount for enhancing an organization’s security posture. By fostering a culture that values and understands the importance of software signing, organizations can ensure the integrity of their software and protect against potential security threats.
@@ 6.4 Contrast to prior work
Distinctions of Refined Model Compared to Existing Frameworks.  While all software supply chain security frameworks agree on the defined threat model, our refinement of this model differs from existing ones in several ways.
@@ 7 Conclusion
Software signing is widely recommended for ensuring software provenance, yet many organizations face challenges when integrating it into their processes. In this first qualitative study, we examine the practices and challenges of software signing through interviews with 18 senior practitioners across 13 organizations. We identified that while software products are signed, the signatures often go unvalidated. Key barriers include infrastructure limitations, lack of expertise, and insufficient resources. Although signing offers strong provenance guarantees, it remains secondary in many organizations’ cybersecurity strategies. Additionally, organizational behaviors are influenced by major incidents like SolarWinds and industry regulations and standards, with concerns raised about the quality and development of these standards.
Our findings shed light on the real-world practices, challenges, and organizational factors shaping software signing adoption, offering guidance for organizations on their journey toward more secure software supply chains.